OpenStudy (anonymous):

when plugging in metric facts, always remember that 1 big unit=# small units. Fill in the fact. ___gram = ___teragram ___nanoliter= ___Liter

OpenStudy (anonymous):

1 teragram=10^12 gram 1 liter= 10^9 nanoliter